Calculus I, Fall 2012 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Chap 04

1. Section 4.2: problems 31, 33, 57, 61, 63, 81.

2. Section 4.3: problems 17, 25, 34, 75, 78, 80.

3. Section 4.4: Problems 23, 29, 35, 41, 58 (Just graph it).

4. Section 4.5: Problems 3, 5, 13, 23, 25, 26, 27(a), 33, 36.

5. Section 4.6: Problems 13, 23(b), 25, 26, 28, 29, 33.

6. Read and compare the difference between Intermediate Value Theorem and Mean Value Theorem.